what does w equal in the equation: 2(v-3)=1-(w+4)
Found 2 solutions by Alan3354, Fombitz:
Answer by Alan3354(69443)

2(v-3)=1-(w+4)
Solve for w
2v-6 = 1 - (w+4)
6-2v = (w+4) - 1
w+4 = 6-2v + 1 = 7-2v
w = 3 - 2v

Answer by Fombitz(32388)